1) A company issued a 5 year, 7% B/P with a face value of $300,000 when the market rate if 8%. How much are the semiannual payments?

2) For 2012, Corn Flake Corporation reported NI of $300,000, interest expense $40,000, and income tax expense $100,000. What is the times interest earned ratio?

3) A company issues an 8 year, 7% B/P with a face value of $200,000 at 98 when the market interest rate is 8%. Interest payments are made semiannually. What is the current market value?

4) A company issued a 5 year, 7% B/P with a face value of $300,000 at 103 when the market rate if 6%. Interest payments are made semiannually. Using the effective interest method, what would be the journal entry for the first payment?

5) Joshua Corporation has current assets $45,000, long-term assets $230,000, current liabilities $36,000, long-term liabilities $100,000, NI $22,000. What is the debt to asset ratio?

6) A company issued a 5 year, 7% B/P with a face value of $300,000 at 103 when the market rate if 6%. Interest payments are made semiannually. Using the straight-line method, what would be the journal entry for the first payment?
Answer #1

1. Semi annual payment each time = 300,000 * 7%* 1/2 = 10500

2. Times interest earned = Income before income taxes and interest expense / Interest expense

= ( 300,000 + 40,000 + 100,000) / 40,000

= 11 times

3. Current market value = 188,347. 70

NPER 16.00 [8*2]
FV 200000
PMT 7000.0 [200,000*14%*1/2]
Rate 4.00%
PV $188,347.70 [-pv(rate,nper,pmt,fv,0)

4.

Account Titles Debit Credit
Interest Expense
[309,000*6%*1/2]
9,270
Premium on bonds payable 1,230
Cash
[$300,000 * 7%*1/2]
                   10,500

5. Debt to assets ratio = Total liabilities / Total assets = (   45,000 + 230,000) / ( 36,000 + 100,000) = 2.022

6.

Account Titles Debit Credit
Interest Expense 9,600
Premium on bonds payable
[9000/10]
900
Cash
[$300,000 * 7%*1/2]
                   10,500